At what temperature will 2.40 moles of chlorine gas exert a pressure of 2.70 atm at a volume of 0.750 L?
1
Expert's answer
2020-05-01T14:40:43-0400

Use the ideal gas law:


PV = nRT


Based on this (R = 0.0821 L·atm/mol·K):


T = PV/nR = (2.70*0.750)/(2.40*0.0821) = 10.28 K


So. at 10.28 K set quantity chlorine gas exert a pressure of 2.70 atm at a volume of 0.750 L